# Image-Text-Image Knowledge Transferring for Lifelong Person Re-Identification with Hybrid Clothing States

### v1

@@ 4.3.2. Results in Unseen Domains
Furthermore, we discuss another important capability of the lifelong ReID model, its generalization ability in unseen domains, given the continuous accumulation of knowledge across different ReID tasks. In Tabs. 4 and 5, we report the results of models trained with Order 1 on the LReID-Hybrid task and trained on the standard LReID benchmarks in unseen domains, respectively.
@@
Similar to the conclusions from the experiments in seen domains, Joint-Train performs better than SFT in the unseen same-cloth scenario but worse in the unseen cloth-changing scenario. The data imbalance in distribution, as well as the amount between datasets, make the model unexpectedly biased toward clothing-related knowledge when different clothing-state datasets are used jointly for training. Additionally, LwF (Li and Hoiem, 2017) and CLIP-ReID (Li et al., 2023) can improve the capability in unseen domains with the help of knowledge distillation and text modality. However, their performance is still far from satisfactory without dealing with knowledge representation mismatch and granularity mismatch properly. As expected, our method outperforms all other competitors on all unseen datasets, significantly showing its superior generalization ability. Thanks to our well-designed modules, $Teata$ can align, transfer and accumulate knowledge in an “image-text-image” closed loop, the consistent and generalized text semantics effectively promote the lifelong learning of ReID models, contributing to superior results in both seen and unseen domains.
Figure 4. Visualization of the learned representations in different domains with various clothing states. For $Teata$, we visualize the representations of (a) each image, and (b) each identity by averaging image representations with the same identity, respectively. We also visualize (c) the learned representations of each identity by SFT for comparison.
@@
Secondly, we train models only on the conventional LReID task, but compare the results on both same-cloth and cloth-changing unseen domains, to explore in depth the generalization ability of lifelong ReID models. As shown in Tab. 5, we observe that (1) Our proposed $Teata$ beats all competitors by a large margin in different scenarios. Although it does not see any samples of changing clothes during the training phase, it is also capable of obtaining impressive results in the cloth-changing scenario. The results demonstrate the consistency and generalization of the text semantics, and that the designs of $Teata$ are effective in not only taking full advantage of text semantics, but also achieving knowledge accumulation in lifelong learning. (2) Since only the same-cloth datasets are used during training in the conventional LReID task, the advanced ReID methods (Ge et al., 2022; Yu et al., 2023) in Tab. 5 achieves inferior results in the unseen cloth-changing domains compare with those in Tabs. 4. (3) For the advanced ReID methods (Ge et al., 2022; Yu et al., 2023), with the performance improvement in the same-cloth scenario, the performance surprisingly decreases in the cloth-changing scenario. We explain that these models may exclusively focus on clothing information due to the overfitting of the training dataset bias, which is not reliable in the cloth-changing scenario. (4) LwF (Li and Hoiem, 2017) and CLIP-ReID (Li et al., 2023) are both built upon the pre-trained vision language model of CLIP. Benefiting from the information complementarity and generalization advantages of image-text multimodality, these two simple baselines show better results than previous lifelong ReID models (Ge et al., 2022; Yu et al., 2023). However, plagued by catastrophic forgetting due to knowledge representation and granularity mismatches, their performance is still sub-optimal. (5) Joint-Train shows no obvious advantage over other previous methods. Additionally, it is worth noting that it may work worse on some same-cloth datasets like CUHK01 (Li et al., 2013) and PRID (Hirzer et al., 2011), and unexpectedly works well in the unseen cloth-changing scenario. One possible reason is the mixture of data from different domains. Joint-Train tries to balance the domain gap, which inevitably sacrifices the performance of datasets with large domain deviations or small scale. On the other hand, clothing appearance in different domains will also appear the representation difference, which encourages the model to explore more robust representations rather than just clothing features. However, it still struggles with learning from different domains, resulting in limited performance.
4.4. Ablation Studies
@@
Table 6. Ablation Studies of our proposed method. The average accuracies of four datasets in Order 1 at the last training step are reported.
@@
+KA-T +SL $\overline{s}_{\rm mAP}^{s}$ $\overline{s}_{\rm R-1}^{s}$ $\overline{s}_{\rm mAP}^{c}$ $\overline{s}_{\rm R-1}^{c}$
@@
Effectiveness of SSP and KAP. We propose SSP and KAP to tackle the knowledge mismatch challenges faced in the LReID-Hybrid task effectively. To handle the challenge of knowledge granularity mismatch, SSP transfers knowledge across the same-cloth and cloth-changing domains with the uniform granularity of text descriptions. As shown in Tab. 6, it effectively achieves performance improvement in both same-cloth and cloth-changing scenarios. Furthermore, after dealing with the knowledge representation mismatch problem via our KA-T and SL designs, our proposed $Teata$ achieves knowledge adaptation effectively, thus showing a significant performance improvement in Tab. 6.
Discussions of our KAP strategy. As discussed in Sec. 3.3, since image and text modalities are aligned in the shared latent space, it would also be effective to use the image representations to achieve knowledge adaptation. We verify it in Tab. 7 (“Ours w/ KA-V”) and observe similar results to ours. Furthermore, we try to remove the text alignment in the first stage and use the unaligned image representations for the classifier initialization (“Ours w/o Align.”). The great performance decrease confirms the effectiveness of the semantic alignment between image and text modalities. It suggests the advantage of leveraging the text space as an intermediary for knowledge transfer and accumulation across domains with hybrid clothing states.
Table 7. Ablation results of knowledge adaptation in our KAP strategy. The average accuracies of four datasets in Order 1 at the last training step are reported.
@@
Visualizations of the learned representations. Our proposed $Teata$ aligns image and text features, thereby transferring the accumulated knowledge in the text space. Benefiting from the consistency, compactness and generalization of text semantics, the learned representations are compact across domains. To demonstrate it, we use t-SNE (Van der Maaten and Hinton, 2008) to visualize the learned representations in different domains. As shown in Fig. 4 (a), the learned image representations are distributed in a unified latent space regardless of domains. The identity representations also exhibit great compactness in Fig. 4 (b). Upon closer observation of Fig. 4 (b), the identity representations on the LTCC (Qian et al., 2020) and PRCC (Yang et al., 2019) datasets are distributed in a similar area. Since both of them involve clothing changes for each pedestrian, their identity representations show robustness to clothing variations, that is exhibit cloth-irrelevant semantics. A similar phenomenon happens on the Market-1501 (Zheng et al., 2015) and MSMT17 (Wei et al., 2018) datasets, where each identity shows consistent appearances in clothing. In contrast, as shown in Fig. 4 (c), the learned identity representations of SFT show great discrepancy for different domains. The learned identity representations on Market-1501 (Zheng et al., 2015), MSMT17 (Wei et al., 2018) and PRCC (Yang et al., 2019) are distributed separately. Due to the great variations in clothing on LTCC (Qian et al., 2020), the identity representations on which are dispersed. The inconsistent semantics can cause knowledge forgetting in the lifelong evolution process. Overall, with the effective text semantics, our proposed $Teata$ achieves the lifelong evolution of ReID models across domains and accumulates abundant knowledge to handle varying clothing states.
5. Conclusion
@@
In this paper, we exploit a more practical task, namely lifelong person re-identification with hybrid clothing states (LReID-Hybrid), which takes a series of same-cloth and cloth-changing domains into account during the lifelong evolution of ReID models. We analyze the challenges of knowledge mismatch across domains with various clothing states and explore the effectiveness of text for LReID-Hybrid. We use CLIP to take advantage of the good alignment between image and text modalities. To complement the text modality, we further propose the Structured Semantic Prompt (SSP) learning and Knowledge Adaptation and Projection (KAP) strategy to tackle the challenges of knowledge mismatches in granularity and representation, respectively. Our proposed novel framework Teata realizes knowledge alignment, transferring and accumulation with an “image-text-image” closed loop. Extensive experimental results on both our proposed LReID-Hybrid and conventional LReID benchmarks demonstrate its effectiveness and superiority. Without any memory for replay, we show that leveraging consistent and generalized text semantics can already promote the performance of LReID significantly. We hope it can inspire more studies to focus on the lifelong evolution of ReID models with hybrid clothing states, promoting the landing of ReID systems.
Appendix
Appendix A More Comparison Results
To better verify the efficacy of our proposed framework Teata, we additionally show results in Orders 2$\sim$6 in Tabs. 8$\sim$12. We compare with one naïve solution of Sequential Fine-Tune (SFT), two rehearsal-free lifelong approaches AKA (Pu et al., 2021) and LwF (Li and Hoiem, 2017), as well as CLIP-ReID (Li et al., 2023) baseline that also uses text to promote ReID learning. Several similar observations can be made from tables that strongly demonstrate the effectiveness and robustness of our proposed method. (1) Teata is consistently superior to AKA, SFT and LwF in different training orders, outperforming them by a large margin. It clearly suggests the advantage of leveraging the text space as an intermediary for knowledge transfer and accumulation across different domains. (2) Although CLIP-ReID achieves relatively better results than other competitors in the same-cloth scenario, it degrades in the cloth-changing scenario. The reason for this phenomenon may be that i) the scale of same-cloth and cloth-changing datasets is unbalanced and ii) the knowledge representations for these two kinds of tasks are mismatched, which eventually leads to catastrophic forgetting, especially the knowledge for cloth-changing ReID. Thanks to our proposed designs of structured semantic prompt (SSP) learning and knowledge adaption and projection module, our Teata could successfully transfer and accumulate knowledge between same-cloth and cloth-changing domains, which bests CLIP-ReID with an overwhelming advantage. (3) As observed and discussed in Sec. 4.3 of our main paper, Joint-Train also performs better in the same-cloth scenario than in the cloth-changing scenario due to the discrepancy in different domains with various clothing states. On the contrary, Teata achieves a good balance in knowledge accumulation across domains, demonstrating the efficacy of our proposed method on lifelong learning.
Table 8. Comparison with the state-of-the-art methods in the LReID-Hybrid setting (Order 2). The results are reported after the last training phase. The best results are shown in bold.
Training Order 2 Market-1501 LTCC PRCC MSMT17 SC Average CC Average
AKA (Pu et al., 2021) 52.2 75.0 7.6 17.6 26.5 25.8 21.8 41.8 37.0 58.4 17.1 21.7
SFT 49.5 74.5 11.6 25.3 22.3 20.8 63.8 84.7 56.7 79.6 22.8 23.1
LwF (Li and Hoiem, 2017) 58.7 79.6 14.2 27.6 32.1 32.0 60.2 82.1 59.5 80.9 23.2 29.8
CLIP-ReID (Li et al., 2023) 56.3 80.0 11.9 26.0 25.2 26.2 67.6 86.5 62.0 83.3 18.9 26.1
Teata (Ours) 82.9 93.0 17.7 35.2 47.4 45.9 65.9 85.0 74.4 89.0 32.6 40.6
Joint-Train 89.1 95.0 13.6 28.8 45.8 47.6 70.3 86.6 79.7 90.8 29.7 38.2
Table 9. Comparison with the state-of-the-art methods in the LReID-Hybrid setting (Order 3). The results are reported after the last training phase. The best results are shown in bold.
Training Order 3 Market-1501 MSMT17 LTCC PRCC SC Average CC Average
AKA (Pu et al., 2021) 58.3 78.6 3.8 10.1 5.6 12.5 33.3 32.7 31.1 44.4 19.5 22.6
SFT 54.8 77.0 37.0 65.7 16.6 35.5 46.9 48.2 45.9 71.4 31.8 41.9
LwF (Li and Hoiem, 2017) 63.9 83.0 49.3 75.3 17.8 31.9 46.9 44.6 56.6 79.2 32.4 38.3
CLIP-ReID (Li et al., 2023) 63.6 82.7 45.9 73.7 18.8 40.1 49.1 48.7 54.8 78.2 34.0 44.4
Teata (Ours) 77.7 90.7 46.4 73.8 23.2 46.9 58.5 57.3 62.1 82.3 40.9 52.1
Joint-Train 89.1 95.0 70.3 86.6 13.6 28.8 45.8 47.6 79.7 90.8 29.7 38.2
Table 10. Comparison with the state-of-the-art methods in the LReID-Hybrid setting (Order 4). The results are reported after the last training phase. The best results are shown in bold.
Training Order 4 LTCC Market-1501 PRCC MSMT17 SC Average CC Average
AKA (Pu et al., 2021) 11.4 25.0 27.9 51.3 29.6 32.6 11.5 27.1 19.7 39.2 20.5 28.8
SFT 11.5 25.8 53.7 78.1 24.2 23.9 63.5 84.4 58.6 81.3 17.9 24.9
LwF (Li and Hoiem, 2017) 13.6 25.0 71.3 87.9 31.7 29.1 55.5 79.4 63.4 83.7 22.7 27.1
CLIP-ReID (Li et al., 2023) 12.0 24.0 61.1 82.5 23.0 23.3 66.2 86.0 63.7 84.3 17.5 23.7
Teata (Ours) 17.4 37.2 72.4 88.9 48.6 49.6 64.9 83.9 68.7 86.4 33.0 43.4
Joint-Train 13.6 28.8 89.1 95.0 45.8 47.6 70.3 86.6 79.7 90.8 29.7 38.2
Table 11. Comparison with the state-of-the-art methods in the LReID-Hybrid setting (Order 5). The results are reported after the last training phase. The best results are shown in bold.
Training Order 5 LTCC Market-1501 MSMT17 PRCC SC Average CC Average
AKA (Pu et al., 2021) 8.3 17.9 20.7 42.3 3.3 9.8 33.1 32.5 12.0 26.1 20.7 25.2
SFT 14.4 29.3 61.2 81.4 44.9 72.1 52.0 54.7 53.1 76.8 33.2 42.0
LwF (Li and Hoiem, 2017) 14.6 28.6 75.8 89.2 48.2 74.3 44.4 41.9 62.0 81.8 29.6 35.3
CLIP-ReID (Li et al., 2023) 14.3 33.2 62.3 81.7 41.7 70.0 42.5 42.3 52.0 75.9 28.4 37.8
Teata (Ours) 20.6 41.6 71.4 87.7 53.8 78.6 61.9 62.9 62.6 83.2 41.3 52.3
Joint-Train 13.6 28.8 89.1 95.0 70.3 86.6 45.8 47.6 79.7 90.8 29.7 38.2
Table 12. Comparison with the state-of-the-art methods in the LReID-Hybrid setting (Order 6). The results are reported after the last training phase. The best results are shown in bold.
Training Order 6 LTCC PRCC Market-1501 MSMT17 SC Average CC Average
AKA (Pu et al., 2021) 11.0 22.4 28.5 30.0 27.6 54.3 10.3 25.5 19.0 39.9 19.8 26.2
SFT 10.6 23.5 22.5 21.1 55.7 79.0 63.6 84.2 59.7 81.6 16.6 22.3
LwF (Li and Hoiem, 2017) 13.4 25.5 29.7 28.2 71.5 86.6 55.4 79.1 63.5 82.9 16.6 26.9
CLIP-ReID (Li et al., 2023) 11.7 23.5 26.2 25.4 61.9 83.2 65.7 85.3 63.8 84.3 19.0 24.5
Teata (Ours) 16.5 36.0 45.4 44.6 75.1 90.0 64.9 84.2 70.0 87.1 31.0 40.3
Joint-Train 13.6 28.8 45.8 47.6 89.1 95.0 70.3 86.6 79.7 90.8 29.7 38.2

### v2

@@ IV-C2 Results in Unseen Domains
We also discuss another important capability of the lifelong ReID model, its generalization ability in unseen domains, given the continuous accumulation of knowledge across different ReID tasks. Without loss of generality, in Tabs. IX and X, we report results of models trained with Order 1 on the LReID-Hybrid task and on the standard LReID benchmarks in unseen domains, respectively.
@@
Similar to the conclusions from experiments in seen domains, Joint-Train performs better than SFT in the unseen same-cloth scenario but worse in the unseen cloth-changing scenario. The imbalance in the distribution and quantity of datasets biases the model towards clothing-related knowledge when different clothing-state datasets are used jointly for training. Additionally, LwF [33] and CLIP-ReID [12] improve capability in unseen domains through knowledge distillation and text modality. However, their performance remains unsatisfactory due to improper handling of knowledge representation and granularity mismatches. Our method, as expected, outperforms all other competitors on all unseen datasets, demonstrating superior generalization ability. Thanks to our well-designed modules, Teata aligns, transfers, and accumulates knowledge in an “image-text-image” closed loop. Consistent and generalized text semantics effectively promote lifelong learning of ReID models, leading to superior results in both seen and unseen domains.
@@
Furthermore, we train models only on the conventional LReID task and compare results on both same-cloth and cloth-changing unseen domains to explore the generalization ability of lifelong ReID models in greater depth. As shown in Tab. X, we observe the following: (1) Teata surpasses all competitors by a large margin in different scenarios. Despite not seeing any samples of changing clothes during training, it achieves impressive results in the cloth-changing scenario. This demonstrates the consistency and generalization of text semantics and the effectiveness of Teata’s designs in utilizing text semantics and achieving knowledge accumulation in lifelong learning. (2) Since only same-cloth datasets are used during training in the conventional LReID task, advanced ReID methods [2, 3] achieve inferior results in unseen cloth-changing domains compared to those in Tabs. IX. (3) Advanced ReID methods [2, 3] show a surprising decrease in performance in the cloth-changing scenario with improvement in the same-cloth scenario. These models may focus exclusively on clothing information due to overfitting on the training dataset bias, which is unreliable in the cloth-changing scenario. (4) LwF [33] and CLIP-ReID [12] leverage the pre-trained vision-language model of CLIP. Benefiting from the complementarity and generalization advantages of image-text multimodality, these baselines perform better than previous lifelong ReID models [2, 3]. However, their performance remains sub-optimal due to catastrophic forgetting caused by knowledge representation and granularity mismatches. (5) Joint-Train shows no obvious advantage over other previous methods and may even perform worse on some same-cloth datasets like CUHK01 [46] and PRID [50], while unexpectedly performing well in the unseen cloth-changing scenario. One possible reason is the mixture of data from different domains. Joint-Train tries to balance the domain gap, sacrificing the performance of datasets with large domain deviations or small scales. On the other hand, representation differences in clothing appearances across domains encourage the model to explore more robust representations beyond just clothing features. However, it still struggles with learning from different domains, resulting in limited performance.
IV-D Ablation Studies
@@
TABLE XI: Ablation studies of our proposed method. The average accuracies of four datasets in Order 1 at the last training step are reported.
@@
KA-T SL $\overline{s}_{\rm mAP}^{s}$ $\overline{s}_{\rm R-1}^{s}$ $\overline{s}_{\rm mAP}^{c}$ $\overline{s}_{\rm R-1}^{c}$
@@
IV-D1 Effectiveness of SSP and KAP
SSP transfers knowledge across the same-cloth and cloth-changing domains with a uniform granularity of text description. As shown in Tab. XI, it results in performance improvement in both same-cloth and cloth-changing scenarios. Furthermore, by addressing the knowledge representation mismatch problem through our KA-T and SL designs, our proposed Teata achieves effective knowledge adaptation, leading to significant performance improvement, as evidenced in Tab. XI. It is worth mentioning that we also observe a similar trend in improving performance for each design in the traditional LReID setting without clothing state variations. SSP can leverage fine-grained attribute semantics in text descriptions by decomposing shared [object] information while keep the discriminative ability using decomposed [content] information. KAP prevents the learning of text embeddings from overfitting to the current task and thus avoids forgetting previous knowledge.
TABLE XII: Ablation results of knowledge adaptation in our KAP strategy. The average accuracies of four datasets in Order 1 at the last training step are reported.
@@ IV-D1 Effectiveness of SSP and KAP
Ours w/o $\mathcal{L}_{proj}$ 64.4 83.5 37.7 46.5
IV-D2 Discussions of the KAP Strategy
As discussed in Sec. III-C, since image and text modalities are aligned in the shared latent space, it would also be effective to use the image representations for knowledge adaptation. We verify this in Tab. XII (“Ours w/ KA-I”) and observe similar results to ours. Furthermore, we remove the text alignment in the first stage and use the unaligned image representations for classifier initialization (“Ours w/o Align.”). The significant performance decrease confirms the effectiveness of semantic alignment between image and text modalities. It suggests the advantage of the text space as an intermediary for knowledge transfer and accumulation across domains. We also ablate the $\mathcal{L}_{proj}$ loss in Tab. XII (“Ours w/o $\mathcal{L}_{proj}$”). The results demonstrate its effectiveness in strengthening the learning of the image encoder by alleviating the loss of text semantics in the knowledge adaptation and projection process.
TABLE XIII: Comparison of one-stage and two-stage training strategies. The average accuracies of four datasets in Order 1 at the last training step are reported.
Methods SC Average CC Average
One Stage 57.8 78.7 34.1 44.6
Two Stages (Ours) 68.4 85.9 40.0 50.2
IV-D3 Necessity of Two-Stage Training
Our proposed two-stage iterative training takes advantage of semantics from the text encoder and shows great superiority in performance to the one-stage baseline SFT and the two-stage training baseline CLIP-ReID in Tabs. $I\simX$. Additionally, we also try a one-stage training variant, which adopts our proposed KAP strategy while uses the contrastive losses $\mathcal{L}_{i2t}$ and $\mathcal{L}_{t2i}$ to achieve our SSP learning at the same time. As shown in Tab. XIII, the variant performs worse, since in the early stage of training the learned text prompts is ineffective in describing pedestrians. The text embeddings with misleading semantics would affect the KAP strategy to optimize the image encoder.
Figure 4: Visualization of the learned representations in different domains with various clothing states. For Teata, we visualize the representations of (a) each image, and (b) each identity by averaging image representations with the same identity, respectively. We also visualize (c) the learned representations of each identity by SFT for comparison.
Figure 5: Influence of different values of $M$. The results are reported after the last training phase in Order 1.
IV-D4 Influence of the Number of Token Pairs
As shown in Fig. 5, a larger value of $M$ generally helps represent different attributes meticulously. We also find that a small number of token pairs can bring good performance. This can be attributed to the fact that the text space is more compact and generalized than the image space. The learnable structured tokens, acting as a container, can distill visual knowledge from images into higher-level representations. When we set $M$ to a large number, such as 32, we observe that it is beneficial in the cloth-changing scenario while providing less gain in the same-cloth scenario. This discrepancy arises from the knowledge granularity mismatch in different clothing states. As discussed in Sec. I, cloth-changing ReID typically requires models to perceive more fine-grained cues in person images, as visual appearance information, which dominates, can be misleading. To achieve a good balance between different knowledge requirements, we set $M=16$ for all our experiments.
IV-D5 Visualizations of the Learned Representations
@@
Our proposed Teata aligns image and text features, transferring the accumulated knowledge into the text space. Benefiting from the consistency, compactness, and generalization of text semantics, the learned representations are compact across domains. To demonstrate this, we use t-SNE [57] to visualize the learned representations in different domains. As shown in Fig. 4 (a), the learned image representations are distributed in a unified latent space regardless of domains. The identity representations also exhibit great compactness in Fig. 4 (b). Upon closer observation of Fig. 4 (b), the identity representations on the LTCC [6] and PRCC [8] datasets are distributed in a similar area. Since both involve clothing changes for each pedestrian, their identity representations show robustness to clothing variations, exhibiting cloth-irrelevant semantics. A similar phenomenon is observed on the Market-1501 [42] and MSMT17 [43] datasets, where each identity shows consistent appearances in clothing. In contrast, as shown in Fig. 4 (c), the learned identity representations of SFT show great discrepancy across different domains. The identity representations on Market-1501[42], MSMT17 [43], and PRCC [8] are distributed separately. Due to the great variations in clothing on LTCC [6], the identity representations are dispersed. These inconsistent semantics can cause knowledge forgetting in the lifelong evolution process. Overall, with the effective use of text semantics, our proposed Teata enables the lifelong evolution of ReID models across domains and accumulates abundant knowledge to handle varying clothing states.
IV-E More Discussions of the LReID-Hybrid Task
We take a further step in LReID by by taking mixed clothing states into account in LReID. For the cloth-changing domains, the newly added dataset may contain persons with different clothing states. For example, for the LTCC [6] training dataset, 46 people have clothing changes while 31 people always wear the same clothing. On the other hand, each person has more than one image, and they are randomly sampled during training. Therefore, there are both cloth-changing and same-cloth images when training in the cloth-changing domains.
To solve LReID-Hybrid task, we first propose SSP learning to transfer knowledge across domains with uniform text description granularity, and then introduce KAP strategy to guide the image representation learning via aligned text embeddings. It is worth noting that our method does not use any information or prior about the clothing states of the datasets.
V Conclusion
@@
Lifelong learning and handling clothing changes are two significant challenges in the application of person ReID. In this paper, we innovatively address both challenges simultaneously by proposing a more practical task: lifelong person re-identification with hybrid clothing states (LReID-Hybrid). This task considers a series of same-cloth and cloth-changing domains during the lifelong evolution of ReID models. As pioneers in this area, we analyze the challenges posed by knowledge mismatches across domains with varying clothing states and explore the effectiveness of text for LReID-Hybrid. We leverage CLIP to take advantage of the robust alignment between image and text modalities. We further propose the Structured Semantic Prompt (SSP) learning and the Knowledge Adaptation and Projection (KAP) strategy to tackle the challenges of knowledge mismatches in granularity and representation, respectively. Our novel framework, Teata, achieves knowledge alignment, transfer, and accumulation within an “image-text-image” closed loop. Extensive experimental results on both LReID-Hybrid and conventional LReID benchmarks demonstrate the superiority of our approach.
We hope our work inspires more studies to focus on the lifelong evolution of ReID models with hybrid clothing states, thus advancing the deployment of ReID systems. Other issues, such as lifelong ReID across visible and infrared modalities, are also significant in real-world applications and remain underexplored. These scenarios present challenges comparable to LReID-Hybrid, and their solutions are far from trivial. Additionally, our work has demonstrated the enhancing role of text in lifelong ReID, which hopes to encourage more research.